# Gait Transitions in Load-Pulling Quadrupeds: Insights from Sled Dogs and a Minimal SLIP Model

### v2

@@ III.2 Hybrid Dynamics of the Quad-load System
We assume that each leg undergoes exactly one touchdown and one liftoff per stride to avoid Zeno behavior [30, 31]. As a result, the eight timing variables $t_{i}$ in (1) partition the hybrid gait cycle into domains j=1:9j=1{:}9, each corresponding to a unique combination of ground-contacting legs. Within each domain, the associated holonomic constraints are enforced using stacked Jacobians $\bm{J}_{j}$ [18], yielding the following differential-algebraic formulation:
@@
where $\Delta$ denotes the impact map that instantaneously resets velocities at transitions. Further implementation details are provided in [7, 32, 18]
@@ IV.1 Observed Distribution of Gait Patterns
Our analysis of high-speed video recordings revealed that sled dogs frequently switch between different galloping gaits during load-pulling. Fig. 4 summarizes the observed distribution of gallops and the frequency of transitions between them for two individual dogs. The arrangement of gait types in the figure follows the same convention as Fig. 2, with transverse gaits (TranL, TranR) placed above and rotary gaits (RotL, RotR) below, and left- and right-leading gaits positioned to the left and right, respectively. The size of each dot is proportional to the number of strides recorded for that gait, while the thickness of the connecting arrows reflects the frequency of transitions between gaits.
@@
Data for Individual 1, compiled from 189 recorded strides, reveals the use of all four galloping gait patterns. Among these, RotL was the most frequently used, occurring in 120 out of 190 strides. This was followed by TranR, which appeared in 50 strides. In contrast, TranL and RotR were used only occasionally. Gait transitions were unevenly distributed. For RotL, 79.83% of consecutive strides remained in the same gait. Transitions to TranR occurred in 15.13% of cases, and to TranL in 5.04%. For strides beginning in TranR, 44.00% remained in TranR, 50.00% transitioned to RotL, and only 6.00% transitioned to TranL. No transitions to RotR were observed from TranR either. These statistics suggest a strong preference for RotL and TranR, while TranL and RotR appear to be infrequent or short-lived. The most common transitions occurred between RotL and TranR (RotL $\leftrightarrow$ TranR), which require changing the touchdown order of only the hindlimbs. In contrast, transitions between RotL and RotR were never observed, likely because they involve a simultaneous reordering of both the forelimb and hindlimb touchdown sequences within a single stride. This more complex coordination change may be less favorable during steady high-speed locomotion.
@@
Data for Individual 2, compiled from 195 recorded strides, reveals a highly consistent gait preference. Nearly all strides were executed in the TranR gait, with only a small fraction performed in RotL. Specifically, TranR accounted for 209 out of 211 strides, while RotL was observed in just a single stride. The other two gaits, TranL and RotR, were not used at all in the recorded dataset. No transitions were observed among three of the four gait types, as TranR and RotL were the only patterns present. For strides beginning in TranR, 99.52% remained in the same gait, while only 0.48%, corresponding to a single stride, transitioned to RotL. Since RotL occurred only once and was not followed by any additional strides of that type, no outbound transitions from RotL could be evaluated. This consistency indicates a strong and stable preference for the TranR gait in Individual 2, with other patterns either avoided or unsustainable. The lack of bidirectional transitions suggests that, in contrast to Individual 1, this dog maintained a highly stable coordination strategy with minimal gait variability across the recorded session.
@@
Figure 4: Distribution and transition dynamics of galloping gaits used by Individual 1 (A) and Individual 2 (B). Each node represents a distinct gait type, with the dot size proportional to its overall usage frequency. Directed arrows indicate transitions between gaits, with arrow thickness representing transition frequency. For each gait node, a corresponding dot plot is shown to illustrate the footfall sequence, using the same color (green for left-leading, blue for right-leading) and background (white for transverse, grey for rotary) convention as defined in Fig. 2. This integrated view links gait prevalence, transitions, and sequence structure in a unified representation.
@@ IV.2 Single Stride Replication
Figure 5: Single-stride solutions for Individual 1, comparing two gaits: (A) TranR and (B) RotL. Left subplots show experimental and simulated footfall sequences; right subplots illustrate experimental and simulated tugline forces along the tugline. Experimental means (solid curves) and variances (shaded regions and error bars) are computed from 50 (TranR) and 120 (RotL) samples.
@@
Due to space constraints, the optimization results for the remaining gaits and for Individual 2 are not visualized here but are included in the publicly available source code and data repository . Table 2 summarizes the $R^{2}$ values across all gait and individual combinations. Footfall timing predictions are highly consistent, with all combinations achieving $R^{2}_{t}=0.99$, indicating precise alignment of simulated and experimental touchdown and liftoff events. Tugline force predictions also show strong agreement, with $R^{2}_{F}$ values ranging from 0.95 to 0.96 across gaits. Stride duration tracking is most accurate for the frequently used RotL and TranR gaits, where $R^{2}_{D}$ exceeds 0.99, and is slightly lower for TranL and RotR, which exhibit more variability. Although TranL and RotR are less frequently used gaits, the SLIP model still maintains reasonable fidelity, with timing and force $R^{2}$ values above 0.93 in all cases. Slightly reduced agreement in TranL, especially for Individual 1, may reflect increased variability in experimental execution or decreased model robustness for rarely used gaits. Nevertheless, the model performs reliably across all conditions tested.
@@ IV.3 Gait Transition Replication
Quantitative analysis supports this interpretation. Defining the stride-to-stride stiffness change as $\Delta k_{\mathrm{swing},\ i}:=k_{\mathrm{swing},\ i}^{n+1}-k_{\mathrm{swing},\ i}^{n}$, we observe substantial and opposing variations in the hindlimbs: $\Delta k_{\mathrm{swing},\ \scriptstyle\text{LH}}=-23.8\,mgl_{0}$ and $\Delta k_{\mathrm{swing},\ \scriptstyle\text{RH}}=+23.1\,mgl_{0}$. In contrast, the forelimbs exhibit only minor adjustments: $\Delta k_{\mathrm{swing},\ \scriptstyle\text{LF}}=-1.38\,mgl_{0}$ and $\Delta k_{\mathrm{swing},\ \scriptstyle\text{RF}}=+1.57\,mgl_{0}$. These values align with the observed swing leg velocity asymmetries and are summarized in Table 3, where the dominant hindlimb contributions are highlighted.
To evaluate whether the identified stiffness pattern is both necessary and sufficient for achieving the transition, we performed a sensitivity analysis by perturbing each $\Delta k_{\text{swing}}$ by $\pm 5\%$ and observing changes in the objective function. As shown in Fig. 6(B, top), the hindlimb curves (LH and RH) show clear minima at zero perturbation, indicating local optimality. In contrast, the forelimb curves (LF and RF) exhibit much smaller variations. The maximum cost increase was 0.54% for LH and 0.39% for RH, compared to only 0.13% and 0.06% for LF and RF, respectively. These results confirm that the transition is highly sensitive to hindlimb stiffness modulation, while the forelimbs remain largely uninvolved.
@@ V Discussion and Conclusion
Gait transitions in quadrupeds are typically associated with changes in locomotor speed. High-speed galloping is often viewed as exhibiting consistent, stereotyped stride patterns [4, 22, 5]. In contrast, our analysis of sprint sled dogs reveals frequent, stride-to-stride transitions among distinct galloping gaits that occur without significant changes in speed, stride duration, or terrain.
This behavior was especially prominent in Individual 1, who alternated between gait types under nearly identical locomotor conditions. This observation brings into question the role of multistability in quadrupedal load pulling. Prior studies have suggested that unloaded quadrupeds may switch lead limbs within an asymmetrical gallop to help balance muscular workload and mitigate fatigue [33]. It has also been suggested that sufficiently large perturbations might induce transitions between distinct galloping modes, though empirical efforts have provided inconclusive evidence [6]. In sustained load-pulling tasks, the ability to switch between mechanically distinct gaits may provide sled dogs a functional advantage for maintaining long-duration performance. Within a team, such transitions could also serve to adaptively redistribute load in response to tugline perturbations, contributing to more robust locomotion. However, the consistent gait observed in Individual 2 suggests that such transitions, if they exist, are not strictly necessary.
@@
Among the parameters examined, swing leg stiffness emerged as a dominant control variable for enabling gait transitions. Biomechanically, this parameter governs how rapidly a leg swings through the air: higher stiffness corresponds to faster oscillation and earlier touchdown. Our results demonstrate that targeted modulation of swing stiffness in just one limb pair was sufficient to replicate the experimentally observed transitions. Specifically, modulating hindlimb stiffness alone reproduced the transition between TranR and RotL, while adjusting forelimb stiffness enabled the transition from RotR to TranR. These adjustments effectively alter the sequencing of leg touchdown events and thus the overall footfall pattern. Sensitivity analyses further confirmed that the cost function was highly responsive to perturbations in the switched limb pair and relatively insensitive to changes in the other limbs. This suggests that sled dogs may exploit stride-wise adjustments in swing timing—achieved through limb-specific stiffness control—as a localized, efficient strategy to switch gaits without global reconfiguration or speed change.
@@
Although our model replicates footfall timings and tugline dynamics with high accuracy, the underlying causes for the observed transitions remain unresolved. One hypothesis is that transitions are influenced by fluctuations in tugline force. However, across 120 RotL strides from Individual 1 and 179 TranR strides from Individual 2, most strides—79% and 99% respectively—did not involve a transition. While some transitions coincided with observable force changes, others occurred without any clear variation, as seen in Fig. 6 and Fig. 7. Due to the small number of transitions, statistical evaluation is not currently feasible. Larger datasets with more frequent gait changes and better-resolved force data would help clarify whether load dynamics play a causal role.
One possible explanation is that variations in pulling force influence the energetic interaction between the dog and the sled, thereby prompting gait transitions. Evaluating this hypothesis is limited by the available data. Among 120 rotary-left (RotL) strides from Individual 1 and 179 transverse-right (TranR) strides from Individual 2, over 79% and 99% of strides remained unchanged, respectively. While changes in tugline force were observed in some transition instances, as shown in Fig. 6 and Fig. 7, similar transitions also occurred with little to no change in force. Again, additional data with more frequent gait switches would be necessary to determine whether load dynamics are a causal factor in gait selection.
Although tri-axial accelerometers were deployed on the same individuals, they were not used in this analysis. Recent work, however, has demonstrated that stance and stride timing can be inferred from inertial measurement unit (IMU) data using machine learning approaches [34], offering a valuable alternative when labeled video is limited or unavailable. To explore this potential, we conducted a preliminary analysis using acceleration data from Individual 1. A bidirectional LSTM model [35] trained on Individual 1’s labeled dataset was able to identify footfall sequences and stance timing with high precision. Across 362 footfalls in a held-out test set, stance events were detected in the correct order, with an average intersection-over-union of 0.869 between predicted and ground truth stance intervals (corresponding to an average timing error of $\sim 10\text{ms}$). A full description of this analysis, including the Python model, training procedure, and evaluation, is provided in . While promising, the temporal precision was insufficient for use in our model fitting. The limited training dataset also precludes broader generalization. Nonetheless, this approach demonstrates a promising path toward accurate, field-deployable gait classification in scenarios where continuous video capture is impractical.
@@
Another limitation of this study lies in the inability to decouple the roles and contributions of individual dogs within the larger team. All experimental data presented here were collected from a ten-dog team hitched in a double-tandem (Alaska) configuration, where the distribution of tensions across the interconnected tether network is not directly measurable. This introduces some ambiguity in interpreting how load dynamics influence gait selection at the individual level. In practice, each dog’s pulling contribution may vary depending on spatial position, terrain conditions, cornering behavior, fatigue, or even team coordination strategies. Whether such factors contribute to observed gait transitions remains an open question.
Future studies could address this limitation by instrumenting each segment of line to measure complete hitch network tension in real time. Additionally, synchronized video and acceleration data from multiple team members could enable modeling of team-level coordination strategies, including how leading versus trailing dogs modulate their gait or force output in response to group dynamics. Decoupling these roles would allow for more precise attribution of observed gait transitions to local mechanical or environmental factors, and could clarify whether transitions are driven by individual control decisions or emergent team-level phenomena.

### v4

@@ III.2 Hybrid Dynamics of the Quad-load System
We assume that each leg undergoes exactly one touchdown and one liftoff per stride to avoid Zeno behavior pace2017piecewise; ames2005sufficient. As a result, the eight timing variables $t_{i}$ in (1) partition the hybrid gait cycle into domains j=1:9j=1{:}9, each corresponding to a unique combination of ground-contacting legs. Within each domain, the associated holonomic constraints are enforced using stacked Jacobians $\bm{J}_{j}$ alqaham2024sixteen, yielding the following differential-algebraic formulation:
@@
where $\Delta$ denotes the impact map that instantaneously resets velocities at transitions. Further implementation details are provided in Ding2024RALSymmetry; westervelt2018feedback; alqaham2024sixteen
@@ IV.1 Observed Distribution of Gait Patterns
Our analysis of high-speed video recordings revealed that sled dogs frequently switch between different galloping gaits during load-pulling. Notably, these gait transitions occurred across relatively narrow speed and stride-frequency ranges. GPS traces from the towed vehicle indicated that both trials were performed at similar forward speeds (Fig. 4). Individual 1 ran at a mean speed of 7.28 m/s (range 6.52–8.38 m/s; CV =5.6=5.6%), while Individual 2 ran at a mean speed of 7.03 m/s (range 6.57–7.69 m/s; CV =2.7=2.7%). Within each trial, speed fluctuated modestly around the mean. Stride frequency likewise remained relatively constant (Individual 1: 2.72 Hz; Individual 2: 2.81 Hz). Transitions occurred even when speed was relatively constant, giving no indication that transitions were restricted to bouts of acceleration or deceleration. To assess stride regularity, we quantified stance and swing durations for each limb. For Individual 1, stance durations were tightly distributed (pooled across limbs: mean =0.0844=0.0844 s, SD =0.0074=0.0074 s, CV $8.8\%$), whereas swing durations were longer and more variable (mean =0.2836=0.2836 s, SD =0.0263=0.0263 s, CV $9.3\%$), consistent with gait transitions being implemented via interlimb phase adjustments during swing (i.e., lead-limb swaps in flight) rather than through delays or irregularities in stance timing. Individual 2 exhibited a similar stance distribution (mean =0.0877=0.0877 s, SD =0.0079=0.0079 s, CV $9.0\%$), while swing durations were slightly shorter and less variable (mean =0.2689=0.2689 s, SD =0.0124=0.0124 s, CV $4.6\%$), consistent with the lower transition rate in this trial.
@@
Fig. 4 also summarizes the observed distribution of gallops and the frequency of transitions between them for two individual dogs. The arrangement of gait types in the figure follows the same convention as Fig. 2, with transverse gaits (TranL, TranR) placed above and rotary gaits (RotL, RotR) below, and left- and right-leading gaits positioned to the left and right, respectively. The size of each dot is proportional to the number of strides recorded for that gait, while the thickness of the connecting arrows reflects the frequency of transitions between gaits.
@@
Data for Individual 1, compiled from 189 recorded strides, reveals the use of all four galloping gait patterns. Among these, RotL was the most frequently used, occurring in 120 out of 189 strides. This was followed by TranR, which appeared in 50 strides. In contrast, TranL and RotR were used only occasionally. Gait transitions were unevenly distributed. For RotL, 79.83% of consecutive strides remained in the same gait. Transitions to TranR occurred in 15.13% of cases, and to TranL in 5.04%. For strides beginning in TranR, 44.00% remained in TranR, 50.00% transitioned to RotL, and only 6.00% transitioned to TranL. No transitions to RotR were observed from TranR either. These statistics suggest a strong preference for RotL and TranR, while TranL and RotR appear to be infrequent or short-lived. The most common transitions occurred between RotL and TranR (RotL $\leftrightarrow$ TranR), which require changing the touchdown order of only the hindlimbs. In contrast, transitions between RotL and RotR were never observed.
@@
Data for Individual 2, compiled from 211 recorded strides, reveals a highly consistent gait preference. Nearly all strides were executed in the TranR gait, with only a small fraction performed in RotL. Specifically, TranR accounted for 209 out of 211 strides, while RotL was observed in just two strides. The other two gaits, TranL and RotR, were not used at all in the recorded dataset. No transitions were observed among three of the four gait types, as TranR and RotL were the only patterns present. For strides beginning in TranR, 99.52% remained in the same gait, while only 0.48%, corresponding to a single stride, transitioned to RotL. For RotL, only two strides were observed, corresponding to one self-transition and one transition to TranR. This consistency indicates a strong and stable preference for the TranR gait in Individual 2, with other patterns either avoided or unsustainable. The lack of bidirectional transitions suggests that, in contrast to Individual 1, this dog maintained a highly stable coordination strategy with minimal gait variability across the recorded session.
@@
Figure 4: Stride frequency, forward speed, and gait transitions for Individual 1 (A,B) and Individual 2 (C,D). (A,C) Time series of stride frequency (top) and forward speed (bottom). Stride frequency was computed as the inverse of the interval between consecutive touchdowns of the trailing hind limb (upper x-axis indicates stride number). Forward speed was estimated from towing-vehicle speed (5 s centered moving average). Vertical grey lines mark identified gait transitions. Gait labels are summarized in Table 2. (B,D) Gait usage and transition network. Each node represents a distinct gait type, with the dot size proportional to its overall usage frequency. Directed arrows indicate transitions between gaits, with arrow thickness representing transition frequency. For each gait node, a corresponding dot plot is shown to illustrate the footfall sequence, using the same color (green for left-leading, blue for right-leading) and background (white for transverse, grey for rotary) convention as defined in Fig. 2. This integrated view links gait prevalence, transitions, and sequence structure in a unified representation. Note that the transition network is based on finite observations (Individual 1: 189 strides with 60 transitions; Individual 2: 211 strides with 2 transitions), and several specific gait types or transition classes are supported by few or zero samples; thus, low-frequency arrows (or missing links) should be interpreted as sample-limited rather than definitive.
[…]
@@ IV.2 Single Stride Replication
Figure 5: Single-stride solutions for Individual 1, comparing two gaits: (A) TranR and (B) RotL. Left subplots show experimental and simulated footfall sequences; right subplots illustrate experimental and simulated tugline forces along the tugline. Experimental means (dashed curves) and variances (shaded regions and error bars) are computed from 50 (TranR) and 120 (RotL) samples.
@@
Due to space constraints, the optimization results for the remaining gaits and for Individual 2 are not visualized here but are included in the publicly available source code and data repository (see the Data and materials availability section). Table 2 summarizes the $R^{2}$ values across all gait and individual combinations. Footfall timing predictions are highly consistent, with all combinations achieving $R^{2}_{t}=0.99$, indicating precise alignment of simulated and experimental touchdown and liftoff events. Tugline force predictions also show strong agreement, with $R^{2}_{F}$ values ranging from 0.95 to 0.96 across gaits. Stride duration tracking is most accurate for the frequently used RotL and TranR gaits, where $R^{2}_{D}$ exceeds 0.99, and is slightly lower for TranL and RotR, which exhibit more variability. Although TranL and RotR are less frequently used gaits, the SLIP model still maintains reasonable fidelity, with timing and force $R^{2}$ values above 0.93 in all cases. Slightly reduced agreement in TranL, especially for Individual 1, may reflect increased variability in experimental execution or decreased model robustness for rarely used gaits. Nevertheless, the model performs reliably across all conditions tested.
@@ IV.3 Gait Transition Replication
Quantitative analysis supports this interpretation. Defining the stride-to-stride stiffness change as $\Delta k_{\mathrm{swing},\ i}:=k_{\mathrm{swing},\ i}^{n+1}-k_{\mathrm{swing},\ i}^{n}$, we observe substantial and opposing variations in the hindlimbs: $\Delta k_{\mathrm{swing},\ \scriptstyle\text{LH}}=-23.8\,mgl_{0}$ and $\Delta k_{\mathrm{swing},\ \scriptstyle\text{RH}}=+23.1\,mgl_{0}$. In contrast, the forelimbs exhibit only minor adjustments: $\Delta k_{\mathrm{swing},\ \scriptstyle\text{LF}}=-1.38\,mgl_{0}$ and $\Delta k_{\mathrm{swing},\ \scriptstyle\text{RF}}=+0.57\,mgl_{0}$. These values align with the observed swing leg velocity asymmetries and are summarized in Table 3, where the dominant hindlimb contributions are highlighted.
To evaluate whether the identified stiffness pattern is both necessary and sufficient for achieving the transition, we performed a sensitivity analysis by perturbing each $\Delta k_{\text{swing}}$ by $\pm 5\%$ independently and observing changes in the objective function. As shown in Fig. 6(D, top), the hindlimb curves (LH and RH) are consistent, in aggregate, with a local minimum at zero perturbation, indicating local optimality with respect to that parameter direction. In contrast, the forelimb curves (LF and RF) exhibit much smaller variations. The maximum cost increase was 0.54% for LH and 0.39% for RH, compared to only 0.13% and 0.05% for LF and RF, respectively. These results confirm that the transition is highly sensitive to hindlimb stiffness modulation, while the forelimbs remain largely uninvolved.
@@ V Discussion and Conclusion
Gait transitions in quadrupeds are typically associated with changes in locomotor speed. High-speed galloping is often viewed as exhibiting consistent, stereotyped stride patterns Hildebrand1977; Hilderbrand1989Quadrupedal; Hoyt1981gaitenergetic. In contrast, our analysis of sprint sled dogs reveals frequent, stride-to-stride transitions among distinct galloping gaits that occur without significant changes in speed, stride duration, or terrain. This interpretation is consistent with symmetry-based theories of locomotion, in which morphology and symmetry structure shape an effective potential landscape with multiple local minima corresponding to distinct gaits Wilshin2017; Golubitsky1999. Although the present quad-load SLIP model is not explicitly formulated as a gradient system, our results are compatible with this perspective. For fixed speed and load, we identify distinct periodic solutions corresponding to different galloping gaits, with sensitivity analyses (Figs.6D and 7D) revealing locally well-defined optima associated with each solution. These optima can be interpreted as coexisting attractors in an effective locomotor landscape shaped by swing-leg stiffness asymmetries. Within this framework, gait transitions may arise from perturbation-driven transitions between neighboring attractors rather than from explicit changes in speed or global control strategy. This behavior was especially prominent in Individual 1, who alternated between gait types under nearly identical locomotor conditions. This observation brings into question the role of multistability in quadrupedal load pulling. Prior studies have suggested that unloaded quadrupeds may switch lead limbs within an asymmetrical gallop to help balance muscular workload and mitigate fatigue walter2007ground. In sustained load-pulling tasks, the ability to switch between mechanically distinct gaits would potentially provide sled dogs a functional advantage for maintaining long-duration performance. Fatigue was not measured here, but both individuals were recorded during comparable training sessions, so we do not consider it an obvious explanation for why Individual 1 switched frequently while Individual 2 remained largely in a single gait.
[…]
@@
Among the parameters examined, swing leg stiffness emerged as a dominant control variable for enabling gait transitions. Biomechanically, this parameter governs how rapidly a leg swings through the air: higher stiffness corresponds to faster oscillation and earlier touchdown. Our results demonstrate that targeted modulation of swing stiffness in just one limb pair was sufficient to replicate the experimentally observed transitions. Specifically, modulating hindlimb stiffness alone reproduced the transition between TranR and RotL, while adjusting forelimb stiffness enabled the transition from RotL to TranL. These adjustments effectively alter the sequencing of leg touchdown events and thus the overall footfall pattern. Sensitivity analyses further confirmed that the cost function was highly responsive to perturbations in the switched limb pair and relatively insensitive to changes in the other limbs. This suggests that sled dogs may exploit stride-wise adjustments in swing timing—achieved through limb-specific stiffness control—as a localized, efficient strategy to switch gaits without global reconfiguration or speed change.
@@
Although our model replicates footfall timings and tugline dynamics with high accuracy, the underlying causes for the observed transitions remain unresolved. One hypothesis is that transitions are influenced by fluctuations in tugline force. However, across 120 RotL strides from Individual 1 and 179 TranR strides from Individual 2, most strides—79% and 99% respectively—did not involve a transition. Importantly, the present dataset is not adequate to quantify the driving mechanisms of gait switching. The number of observed transitions is too small to estimate load- or speed-dependent transition probabilities with confidence. Additional data is necessary to test whether and how mechanical perturbations bias gait switching during high-speed load pulling.
In interpreting the transition fits and sensitivity results, several limitations should be kept in mind. First, our sensitivity analysis is centered on transition strides rather than mapping cost gradients around canonical TranR or RotL solutions, in part because the parameterization is effectively high-dimensional and over-actuated, so multiple parameter directions can yield acceptable fits and control-like gradients may not be unique. Secondly, we used a model in which the primary dissipation arises from load friction and did not include explicit joint-level damping. Because damping can strongly influence task-level stability Heim2020 and may be especially consequential for transition strides, our conclusions regarding transition robustness and stability should be interpreted within this simplified structure. Third, in Fig. 6, the cost is evaluated on a single two-stride transition instance, so we cannot strictly rule out instance-specific overfitting, even though similar qualitative stiffness-change structure appears across multiple realizations of the same transition type. More broadly, we do not claim that alternative parameters have been eliminated as contributors, and because the optimization is gradient-based we have not systematically characterized sensitivity to initial guesses or the structure of the optimization landscape. Finally, stiffness modulation is one of the primary available levers to reorganize timing in our reduced-order model. While the results support a consistent model-based relationship between differential swing stiffness and footfall reordering, establishing biological causality would require predictive perturbations or independent validation.
Future work should use protocols that prescribe speed/acceleration profiles to quantify how transition probability depends on these factors. Although we restricted our analysis to straight, steady-running segments, speed was not experimentally controlled. Treadmill experiments, however, could enforce constant speeds, or—if the animal remains harnessed on the treadmill—could also apply controlled perturbations to quantify gait-switching responses. Additionally, the short duration and moderate intensity of the bouts analyzed here preclude strong inference about fatigue-driven switching. Future experiments will require longer recordings and fatigue proxies (e.g., heart rate, metabolic estimates, or limb-loading asymmetries) to quantify whether lead-switching frequency increases with cumulative workload under load pulling. In parallel, model extensions that incorporate limb-specific costs or fatigue-like dynamics could help interpret such experiments and generate testable predictions for when gait switching should emerge under load, including whether within-family switches correspond to larger inferred costs than transverse–rotary switches.
@@
Another limitation of this study lies in the inability to decouple the roles and contributions of individual dogs within the larger team. All experimental data presented here were collected from a ten-dog team hitched in a double-tandem (Alaska) configuration, where the distribution of tensions across the interconnected tether network is not directly measurable. This introduces some ambiguity in interpreting how load dynamics influence gait selection at the individual level. In practice, each dog’s pulling contribution may vary depending on spatial position, terrain conditions, cornering behavior, fatigue, or even team coordination strategies. Whether such factors contribute to observed gait transitions remains an open question. Future studies could address this limitation by instrumenting each segment of line to measure complete hitch network tension in real time. Additionally, synchronized video and acceleration data from multiple team members could enable modeling of team-level coordination strategies, including how leading versus trailing dogs modulate their gait or force output in response to group dynamics. Decoupling these roles would allow for more precise attribution of observed gait transitions to local mechanical or environmental factors, and could clarify whether transitions are driven by individual control decisions or emergent team-level phenomena.
A primary barrier to addressing the limitations above is data scale. Rigorous transition statistics will require substantially more annotated strides across individuals and conditions. Although treadmill protocols can prescribe speed and perturbations in a controlled setting, field studies remain challenging both to deploy at scale and to annotate. Recent work, however, has demonstrated that stance and stride timing can be inferred from inertial measurement unit (IMU) data using machine learning approaches serra2020improving, offering a valuable alternative when labeled video is limited or unavailable. To assess feasibility in our setting, we trained a bidirectional LSTM sak2014long on Individual 1 using stance labels annotated from the synchronized 240 fps video and discretized onto the 120 Hz IMU time base. In a held-out test set (362 footfalls), the model recovered the correct footfall order and achieved an average intersection-over-union (IoU) of 0.869 for stance intervals. Typical discrepancies were on the order of one IMU timestep ($\sim$10 ms), near the effective temporal resolution of the sensor. Code and evaluation details are provided in our repository (Sec. Data Availability). While promising, the current training set is still too limited for broader generalization. Nonetheless, IMU-only gait inference could substantially reduce reliance on continuous video and enable the larger datasets needed to quantify speed- and load-dependent transition probabilities.
Implications
Together, our stride-resolved measurements and SLIP modeling framework provide a new window into high-speed quadrupedal locomotion under sustained external loading. Our empirical results provide an additional example of discrete gait switching consistent with quadrupedal multistability, here observed at high speed during sustained load pulling. Although we do not identify a specific driver of switching, sprint sled dogs alternate between transverse and rotary galloping patterns within a comparatively narrow speed range, suggesting that gait choice is not a deterministic function of speed alone. It should be acknowledged that transverse–rotary switching corresponds to a relatively small reconfiguration of interlimb phasing compared with classical trot–gallop transitions Nevertheless, transitions among closely related galloping solutions are still consistent with multistable locomotor dynamics in a regime where limb timing is otherwise highly constrained, and may support robust locomotion under perturbation while redistributing muscular demand over time. Our SLIP-based model reproduces observed switching using targeted modulation of a small subset of physically interpretable parameters, including stiffness asymmetries and swing-phase adjustments. This suggests that low-dimensional control knobs may be sufficient to organize gait switching in quadruped–load systems and providing a principled starting point for future, load-aware gait-switching controllers.
Animal Subjects and Ethics
The data analyzed in this study were collected from two sprint-racing sled dogs (one male, one female; 1.5 and 2.5 years old, respectively) from a professional sprint-racing kennel. Both dogs were bred and trained for the sport of sprint racing and are representative of the Eurohound type. For the run analyzed here, dog positions within the team were chosen by their owner based on each dog’s temperament and aptitude. All procedures were conducted in accordance with institutional and national guidelines for research involving animals. Protocols were reviewed and approved by the Institutional Animal Care and Use Committee at Georgia Institute of Technology (protocol BHAMLA-A100575U-10/06/2025). Written informed consent was obtained from the dogs’ owners prior to participation. Dogs were housed and cared for by their owners according to standard husbandry practices. All experimental sessions were integrated into the dogs’ normal training routine and were discontinued immediately if any signs of distress were observed.